# Overlap Gap and Computational Thresholds in the Square Wave Perceptron

### v2

@@ 4.2.2 The $\delta\neq 0$ Case
When $\delta$ is different from zero, the OGP transition lines retain aspects similar to those examined in the storage setting. In particular, for $\delta$ sufficiently large, the annealed approximation gives estimates of $\alpha_{\mathrm{OGP}}$ that are above $\alpha_{T}$, see Fig. 6(a). This is clearly wrong, as it has already been shown in the ABP [16] that one can find a generic solution only strictly below $\alpha_{T}$. Using an RS ansatz solves this problem (see Fig. 6(b)), and in particular for $\delta\to\infty$ and $m=4$ one obtains $\alpha^{\mathrm{rs}}_{\mathrm{OGP}}\simeq 1.15$, again similar to the value reported in [16]. However, as in the storage case, the RS ansatz is not enough to correctly compute the value of $\alpha_{\mathrm{OGP}}$ for large $m$.
@@ 4.3 Recovery of the Planted Signal
In this section we address the recovery problem, which consists in the inference of the planted solution. It can be useful to distinguish between weak recovery, which corresponds to inferring a finite fraction of bits of the teacher, from full recovery, which corresponds to inferring all the bits. Full recovery below $\alpha_{T}(\varphi)$ is impossible, since there is an exponential number of solutions and the teacher is indistinguishable from them. Above $\alpha_{T}(\varphi)$, full recovery is instead information-theoretically possible. Indeed, since the planted solution is the unique solution, an exhaustive search would find it. Interestingly, depending on the activation function and the domain of definition of the weights, it might happen that the best efficient algorithms require a fraction of patterns strictly larger than $\alpha_{T}(\varphi)$ in order to fully recover the teacher; namely they display a statistical-to-computational gap. This is the case, for example, both in planted $\text{SBP}_{\kappa}$ and in ABP [16, 52]. From the static point of view, in hard inference problems, there is a region above $\alpha_{T}(\varphi)$ characterized by a coexistence between a thermodynamically stable phase, described by the fixed point of the replica-symmetric free energy associated with the planted solution, and a so-called metastable phase, associated with another locally stable fixed point with a lower free entropy. The spinodal transition $\alpha_{r}$ is defined as the value of $\alpha$ s.t. for $\alpha>\alpha_{r}$ the metastable fixed point becomes locally unstable. In general, no known efficient algorithm can recover the planted signal below $\alpha_{r}$. An exception is represented by the peculiar case of planted $k$-XOR-SAT, where the metastable phase extends all the way up to $\alpha=O(N^{k/2})$ [19], but due to its linear structure, the problem can be solved in polynomial time by Gaussian elimination. In “hard” inference problems, the spinodal transition is believed to represent a fundamental barrier to efficient algorithms. In particular, belief propagation (BP) and variations of spectral methods saturate the spinodal threshold [57]. Numerical evidences support the same for a replicated version of Monte Carlo, called replicated simulated annealing [58]. In binary perceptron problems, above the spinodal transition, BP and Approximate Message Passing (AMP) (which for these problems is equivalent to BP in the large $N$ limit) converges to the planted solution. Given a realization of the dataset, in AMP one writes down iterative equations for single-variable marginals of the uniform measure over all solutions of the perceptron (see G). Since we are considering the case of binary weights $w_{i}$, each marginal $\mu_{i}(w_{i})$ can be parameterized by a single number, e.g. the magnetization $m_{i}=\sum_{i}\mu_{i}(w_{i})\,w_{i}$. The $m_{i}$ depends on the site $i$. It turns out that this dependence can be characterized statistically in the large $N$ limit. Indeed, it is possible to write closed equations (state evolution) for the probability distributions of the magnetizations iterated by AMP. A remarkable fact is that the fixed points of the state evolution equations are in correspondence with the fixed points of the replica symmetric free entropy. For this reason, the properties of the metastable fixed point that one finds from the analysis of the RS free entropy is crucial for the behavior of AMP. Indeed, as we are going to see, an initialization of the AMP messages at zero overlap with the planted signal (non-informed initialization), leads in the hard phase to the metastable fixed point. Note that AMP returns marginals, not configurations of the hypercube. However, in perceptrons the fixed point associated with the planted signal has overlap $p=1$ with the teacher. This is a consequence of the fact that the teacher is an isolated solution. In other words, on this fixed point the magnetizations are all polarized to ±$\pm 1$, namely they correspond to a configuration of the hypercube (the teacher), and therefore AMP is an actual solver for the recovery problem.
@@
Let us discuss how to compute the spinodal transition. We call $\phi(p,\alpha)$ the replica symmetric free entropy on the Nishimori line, where $p=q$ (see B). In order to identify $\alpha_{r}$, one should find $p^{*}<1$ such that:
@@
The first equation asks $p^{*}$ to be a fixed point. The second is the marginality requirement of the fixed point that identifies $\alpha_{r}$. Equations (42) can be solved numerically. In the case of the SWP, they lead to the predictions shown in Fig. 6(a).
@@
If the activation function is balanced, namely if
@@
then $H_{0}(\tau)=1/2$, and therefore $g_{0}=-\log{2}$. The previous formulas tell us that in order for the metastable fixed point to be non-informative one should have $H_{1}(\tau)=0$ (because one should have $g_{1}=0$). If the activation is even, this condition is automatically satisfied. If $g_{1}=0$, then the non-informative fixed point is stable up to $\alpha_{r}=1/g_{2}$. If $g_{2}=0$, then the non-informative point is always stable for $\alpha=O(1)$, meaning that AMP cannot do better than random guessing for $\alpha=O(1)$. For example, in the case of the SWP, for $\delta\rightarrow 0$, one has that $g_{1},g_{2}\rightarrow 0$. In the next section, we discuss a simple modification of the SWP that reproduces the same phenomenology, without having to take an infinite frequency limit.
@@ 4.3.1 The Spinodal Transition in the Reversed Wedge Perceptron
The activation (48) is known as the reversed wedge perceptron (RWP) [59, 60, 61, 62, 63]. Since (48) is odd, we automatically have $H_{2}(\tau)=0$. Putting $H_{1}=0$, one finds $\gamma_{\star}=\sqrt{2\log{2}}$. Let us call $\varepsilon=|\gamma-\gamma_{\star}|$. The following scalings are found for small $\varepsilon$:
$p^{*}\approx P\,\varepsilon,\quad\alpha_{r}\approx A\,\varepsilon^{-1}\,,$ (49)
therefore in the limit $\gamma\rightarrow\gamma_{\star}$ the metastable state does not destabilize for $\alpha=O(1)$ within a RS ansatz. Using the scalings (49) in equations (42), where the expression for the free entropy is given in B), and expanding for small $\varepsilon$, it is possible to compute $P,A$ defined in Eq. (49). In order to fix $P$ and $A$, one has to expand the energetic term up to $p^{3}$. One finds:
@@
At this point using ansatz (49), we get a closed equation for $P,A$:
$\begin{cases}-P+2A\left(a_{1}(\varepsilon)+3\,a_{3}(\varepsilon)\,P^{2}\right)=0\,,\\ -1+12\,A\,a_{3}(\varepsilon)\,P=0\,.\end{cases}$ (53)
Note that the term proportional to $a_{2}$ does not contribute since it is $O(\varepsilon^{4})$. Therefore one finds,
@@
In Fig. 7 we show a comparison between the scalings of Eq. (54) and the numerical solution of Eq. (42) for small $\varepsilon$. In Figs. 8(a) and 8(b) we compare numerical experiments performed with AMP with the analytical prediction for the spinodal transition for different values of $\gamma$.
@@
Figure 8: (a): Success rate of AMP in finding the planted signal when using the RWP activation function in (48). We plot the success rate as a function of $\gamma$, for different values of $\alpha$. Each point is averaged over $20$ independent samples of size $N=5\times 10^{3}$. (b): Hard-Easy transition $\alpha_{r}$ as a function of $\gamma$. The transition diverges at $\gamma=\gamma_{\star}=\sqrt{2\log{2}}$. The continuous line is computed by solving numerically the RS saddle-point equations. The points are numerical simulations of AMP obtained by averaging over $20$ independent samples of size $N=5\times 10^{3}$.
@@ 4.3.2 Bias Detection in the Planted Ensemble
$x^{(y)}_{k}=2^{P}\,\mathds{E}_{\{\boldsymbol{x^{\mu}}\}}\left[\prod_{\mu}\Theta\left(y\,\varphi\left(\frac{1}{\sqrt{N}}\sum_{i}x^{\mu}_{i}w_{i}^{T}\right)\right)\frac{1}{P}\sum_{\mu}x^{\mu}_{k}\right]\,,$ (56)
@@
$x^{(y)}_{k}=\frac{2}{P}\sum_{\mu}\,\mathds{E}_{\{\boldsymbol{x^{\mu}}\}}\left[\Theta\left(y\,\varphi\left(\frac{1}{\sqrt{N}}\sum_{i}x^{\mu}_{i}w_{i}^{T}\right)\right)\,x^{\mu}_{k}\right]=2\int D_{\Sigma}(u,v)\,\Theta\left(y\,u\right)\,v\,,$ (57)
@@
$\Sigma\approx\begin{pmatrix}1&w_{k}^{T}/\sqrt{N}\\ w_{k}^{T}/\sqrt{N}&1\end{pmatrix}\,.$ (58)
@@
The previous formulas tell us that the barycenter is generally biased towards the teacher. If one chooses the activation s.t. $H_{1}=0$ the barycenter becomes non-informative, and also the first term $g_{1}$ of the expansion of the free energy vanishes, implying that the point at zero overlap with the teacher in the state evolution equations becomes a fixed point.
@@
$\mathds{E}_{\boldsymbol{x}^{\mu}}x^{\mu}_{i}\,x^{\mu}_{j}\Theta\left(y^{\mu}\varphi\left(\frac{1}{\sqrt{N}}\sum_{i}x^{\mu}_{i}w^{T}_{i}\right)\right)=\\ =\int\frac{d\hat{h}^{\mu}dh^{\mu}}{2\pi}\mathrm{e}^{i\hat{h}^{\mu}h^{\mu}}\Theta\left(y^{\mu}\varphi(h^{\mu})\right)\left[\prod_{k\neq i,j}\mathds{E}_{x^{\mu}_{k}}\mathrm{e}^{-i\hat{h}^{\mu}\frac{x^{\mu}_{k}w^{T}_{k}}{\sqrt{N}}}\right]\left[\mathds{E}_{x^{\mu}_{i}}x^{\mu}_{i}\mathrm{e}^{-i\hat{h}^{\mu}\frac{x^{\mu}_{i}w^{T}_{i}}{\sqrt{N}}}\right]\left[\mathds{E}_{x^{\mu}_{j}}x^{\mu}_{j}\mathrm{e}^{-i\hat{h}^{\mu}\frac{x^{\mu}_{j}w^{T}_{j}}{\sqrt{N}}}\right]=\\ =\frac{w_{i}^{T}w_{j}^{T}}{N}\int Dh\,(h^{2}-1)\,\Theta\left(y^{\mu}\varphi(h)\right)=\frac{w_{i}^{T}w_{j}^{T}}{N}H_{2}(y^{\mu})\,,$ (61)
@@ A.1 Annealed Computation
$\begin{split}\mathds{E}\mathcal{N}_{m}(q)&=\int\prod_{a}dp_{a}\,\left(\int D_{\Sigma}\boldsymbol{u}\prod_{a=1}^{m}\Theta(\varphi(u_{0})\varphi(u_{a}))\right)^{P}\,\frac{1}{2^{N}}\sum_{w_{0}\cdots w_{m}}\,\prod_{a<b}\delta\left(q-\frac{\boldsymbol{w}_{a}\cdot\boldsymbol{w}_{b}}{N}\right)\,\prod_{a}\delta\left(p_{a}-\frac{\boldsymbol{w}_{0}\cdot\boldsymbol{w}_{a}}{N}\right)=\\ &=\int\prod_{a}dp_{a}\,\mathrm{e}^{mN\,\phi^{\text{ann}}_{m}(q,p_{1},\dots,p_{m})}\,,\end{split}$ (66)
@@
$G_{S}^{(m)}(q,p_{1},\dots,p_{m})=\lim_{N\rightarrow\infty}\frac{1}{mN}\log{\frac{1}{2^{N}}\sum_{\boldsymbol{w}_{0}\cdots\boldsymbol{w}_{m}}\prod_{a<b}\delta\left(q-\frac{\boldsymbol{w}_{a}\cdot\boldsymbol{w}_{b}}{N}\right)\,\prod_{a}\delta\left(p_{a}-\frac{\boldsymbol{w}_{0}\cdot\boldsymbol{w}_{a}}{N}\right)}\,,$ (68)
@@
$\frac{1}{2^{N}}\sum_{\boldsymbol{w}_{0}\cdots\boldsymbol{w}_{m}}\prod_{a<b}\delta\left(q-\frac{\boldsymbol{w}_{a}\cdot\boldsymbol{w}_{b}}{N}\right)\,\prod_{a}\delta\left(p_{a}-\frac{\boldsymbol{w}_{0}\cdot\boldsymbol{w}_{a}}{N}\right)=\int\prod_{a}\frac{d\hat{p}_{a}}{2\pi/N}\prod_{a<b}\frac{d\hat{q}_{b}}{2\pi/N}\,\mathrm{e}^{-Nq\sum_{a<b}\hat{q}_{ab}-Np\sum_{a}\hat{p}_{a}}\times\\ \times\left(\frac{1}{2}\sum_{w_{0}\cdots w_{m}}\mathrm{e}^{\sum_{a<b}\hat{q}_{ab}w_{a}w_{b}+w_{0}\sum_{a}\hat{p}_{a}w_{a}}\right)^{N}\,.$ (71)
@@
The last identity in Eq. 75 follows noting that,
@@
$-\frac{1}{2}\sum_{ab}q^{ab}\hat{h}^{a}\hat{h}^{b}=-\frac{q}{2}\left(\sum_{a}\hat{h}^{a}\right)^{2}-\frac{1-q}{2}\sum_{a}(\hat{h}^{a})^{2}\,.$ (79)

### v4

@@ 4.2.2 The $\delta\neq 0$ Case
When $\delta$ is different from zero, the OGP transition lines retain aspects similar to those examined in the storage setting. In particular, for $\delta$ sufficiently large, the annealed approximation gives estimates of $\alpha_{\mathrm{OGP}}$ that are above $\alpha_{T}$, see Fig.˜6(a). This is clearly wrong, as it has already been shown in the ABP [16] that one can find a generic solution only strictly below $\alpha_{T}$. Using an RS ansatz solves this problem (see Fig.˜6(b)), and in particular for $\delta\to\infty$ and $m=4$ one obtains $\alpha^{\mathrm{rs}}_{\mathrm{OGP}}\simeq 1.15$, again similar to the value reported in [16]. However, as in the storage case, the RS ansatz is not enough to correctly compute the value of $\alpha_{\mathrm{OGP}}$ for large $m$.
@@ 4.3 Recovery of the Planted Signal
In this section we address the recovery problem, which consists in the inference of the planted solution. It can be useful to distinguish between weak recovery, which corresponds to inferring a finite fraction of bits of the teacher, from full recovery, which corresponds to inferring all the bits. Full recovery below $\alpha_{T}(\varphi)$ is impossible, since there is an exponential number of solutions and the teacher is indistinguishable from them. Above $\alpha_{T}(\varphi)$, full recovery is instead information-theoretically possible. Indeed, since the planted solution is the unique solution, an exhaustive search would find it. Interestingly, depending on the activation function and the domain of definition of the weights, it might happen that the best efficient algorithms require a fraction of patterns strictly larger than $\alpha_{T}(\varphi)$ in order to fully recover the teacher; namely they display a statistical-to-computational gap. This is the case, for example, both in planted $\text{SBP}_{\kappa}$ and in ABP [16, 56]. From the static point of view, in hard inference problems, there is a region above $\alpha_{T}(\varphi)$ characterized by a coexistence between a thermodynamically stable phase, described by the fixed point of the replica-symmetric free entropy associated with the planted solution, and a so-called metastable phase, associated with another locally stable fixed point with a lower free entropy. The spinodal transition $\alpha_{r}$ is defined as the value of $\alpha$ s.t. for $\alpha>\alpha_{r}$ the metastable fixed point becomes locally unstable. In general, no known efficient algorithm can recover the planted signal below $\alpha_{r}$. An exception is represented by the peculiar case of planted $k$-XOR-SAT, where the metastable phase extends all the way up to $\alpha=O(N^{k/2})$ [22], but due to its linear structure, the problem can be solved in polynomial time by Gaussian elimination. In “hard” inference problems, the spinodal transition is believed to represent a fundamental barrier to efficient algorithms. In particular, belief propagation (BP) and variations of spectral methods saturate the spinodal threshold [61]. Numerical evidences support the same for a replicated version of Monte Carlo, called replicated simulated annealing [62]. In binary perceptron problems, above the spinodal transition, BP and Approximate Message Passing (AMP) (which for these problems is equivalent to BP in the large $N$ limit) converges to the planted solution. Given a realization of the dataset, in AMP one writes down iterative equations for single-variable marginals of the uniform measure over all solutions of the perceptron (see I). Since we are considering the case of binary weights $w_{i}$, each marginal $\mu_{i}(w_{i})$ can be parameterized by a single number, e.g. the magnetization $m_{i}=\sum_{i}\mu_{i}(w_{i})\,w_{i}$. The $m_{i}$ depends on the site $i$. It turns out that this dependence can be characterized statistically in the large $N$ limit. Indeed, it is possible to write closed equations (state evolution) for the probability distributions of the magnetizations iterated by AMP. A remarkable fact is that the fixed points of the state evolution equations are in correspondence with the fixed points of the replica symmetric free entropy. For this reason, the properties of the metastable fixed point that one finds from the analysis of the RS free entropy is crucial for the behavior of AMP. Indeed, as we are going to see, an initialization of the AMP messages at zero overlap with the planted signal (non-informed initialization), leads in the hard phase to the metastable fixed point. Note that AMP returns marginals, not configurations of the hypercube. However, in perceptrons the fixed point associated with the planted signal has overlap $p=1$ with the teacher. This is a consequence of the fact that the teacher is an isolated solution. In other words, on this fixed point the magnetizations are all polarized to $\pm 1$, namely they correspond to a configuration of the hypercube (the teacher), and therefore AMP is an actual solver for the recovery problem.
@@
Let us discuss how to compute the spinodal transition. We call $\phi(p,\alpha)$ the replica symmetric free entropy on the Nishimori line, where $p=q$ (see C). In order to identify $\alpha_{r}$, one should find $p^{*}<1$ such that:
@@
The first equation asks $p^{*}$ to be a fixed point. The second is the marginality requirement of the fixed point that identifies $\alpha_{r}$. Equations (44) can be solved numerically. In the case of the SWP, they lead to the predictions shown in Fig.˜6(a).
@@
are the first Hermite polynomials. See, e.g., Refs. [63, 64, 65, 66] for representative works analyzing the link between weak recovery by efficient algorithms and the Hermite decomposition of the activation.
If the activation function is balanced, that is, if
@@
then $H_{0}(\tau)=1/2$, and therefore $g_{0}=-\log{2}$. The previous formulas tell us that in order for the metastable fixed point to be non-informative one should have $H_{1}(\tau)=0$ (because one should have $g_{1}=0$). If the activation is even, this condition is automatically satisfied. If $g_{1}=0$, then the non-informative fixed point is stable up to $\alpha_{r}=1/g_{2}$. If $g_{2}=0$, then the non-informative point is always stable for $\alpha=O(1)$, meaning that AMP cannot do better than random guessing for $\alpha=O(1)$. For example, in the case of the SWP, for $\delta\rightarrow 0$, we have $g_{1},g_{2}\rightarrow 0$.
In what follows, we give a physical interpretation of the Hermite decomposition and show that a simple modification of the SWP reproduces the divergence of the AMP recovery threshold without requiring to take an infinite-frequency limit.
@@ 4.3.1 The Spinodal Transition in the Reversed Wedge Perceptron
The activation (50) is known as the reversed wedge perceptron (RWP) [67, 68, 69, 70]. Since (50) is odd, we automatically have $H_{2}(\tau)=0$. Putting $H_{1}=0$, one finds $\gamma_{\star}=\sqrt{2\log{2}}$. Let us call $\varepsilon=|\gamma-\gamma_{\star}|$. The following scalings are found for small $\varepsilon$:
$p^{*}\approx U\,\varepsilon,\quad\alpha_{r}\approx A\,\varepsilon^{-1}\,$, (51)
therefore, in the limit $\gamma\rightarrow\gamma_{\star}$ the metastable state does not destabilize for $\alpha=O(1)$ within a RS ansatz. Using the scalings (51) in equations (44), where the expression for the free entropy is given in C, and expanding for small $\varepsilon$, it is possible to compute $U,A$ defined in Eq. (51). In order to fix $U$ and $A$, one has to expand the energetic term up to $p^{3}$. One finds:
@@
At this point using ansatz (51), we get a closed equation for $U,A$:
$\begin{cases}-U+2A\left(a_{1}(\varepsilon)+3\,a_{3}(\varepsilon)\,U^{2}\right)=0\,,\\ -1+12\,A\,a_{3}(\varepsilon)\,U=0\,.\end{cases}$ (55)
Note that the term proportional to $a_{2}$ does not contribute since it is $O(\varepsilon^{4})$. Therefore, one finds:
@@
In Fig.˜7 we show a comparison between the scalings of Eq. (56) and the numerical solution of Eq. (44) for small $\varepsilon$. In Figs.˜8(a) and 8(b) we compare numerical experiments performed with AMP with the analytical prediction for the spinodal transition for different values of $\gamma$.
@@
Figure 8: (a): Success rate of AMP in finding the planted signal when using the RWP activation function in (50). We plot the success rate as a function of $\gamma$, for different values of $\alpha$. Each point is averaged over 2020 independent samples of size $N=5\times 10^{3}$. (b): Hard-Easy transition $\alpha_{r}$ as a function of $\gamma$. The transition diverges at $\gamma=\gamma_{\star}=\sqrt{2\log{2}}$. The continuous line is computed by solving numerically the RS saddle-point equations. The points are numerical simulations of AMP obtained by averaging over 2020 independent samples of size $N=5\times 10^{3}$.
@@ 4.3.2 Bias Detection in the Planted Ensemble
$x^{(y)}_{k}=2^{P}\,\mathds{E}_{\{\boldsymbol{x^{\mu}}\}}\left[\prod_{\mu}\Theta\left(y\,\varphi\left(\frac{1}{\sqrt{N}}\sum_{i}x^{\mu}_{i}w_{i}^{*}\right)\right)\frac{1}{P}\sum_{\mu}x^{\mu}_{k}\right]\,$, (58)
@@
$x^{(y)}_{k}=\frac{2}{P}\sum_{\mu}\,\mathds{E}_{\{\boldsymbol{x^{\mu}}\}}\left[\Theta\left(y\,\varphi\left(\frac{1}{\sqrt{N}}\sum_{i}x^{\mu}_{i}w_{i}^{*}\right)\right)\,x^{\mu}_{k}\right]=2\int D_{\Sigma}(u,v)\,\Theta\left(y\,\varphi(u)\right)\,v\,$, (59)
@@
$\Sigma=\begin{pmatrix}1&w_{k}^{*}/\sqrt{N}\\ w_{k}^{*}/\sqrt{N}&1\end{pmatrix}\,$. (60)
@@
The previous formulas tell us that the barycenter is generally biased towards the teacher. If one chooses the activation s.t. $H_{1}=0$ the barycenter becomes non-informative, and also the first term $g_{1}$ of the expansion of the free entropy vanishes, implying that the point at zero overlap with the teacher in the state evolution equations becomes a fixed point.
@@
$\mathds{E}_{\boldsymbol{x}^{\mu}}x^{\mu}_{i}\,x^{\mu}_{j}\Theta\left(y^{\mu}\varphi\left(\frac{1}{\sqrt{N}}\sum_{i}x^{\mu}_{i}w^{*}_{i}\right)\right)=\\ =\int\frac{d\hat{h}^{\mu}dh^{\mu}}{2\pi}\mathrm{e}^{i\hat{h}^{\mu}h^{\mu}}\Theta\left(y^{\mu}\varphi(h^{\mu})\right)\left[\prod_{k\neq i,j}\mathds{E}_{x^{\mu}_{k}}\mathrm{e}^{-i\hat{h}^{\mu}\frac{x^{\mu}_{k}w^{*}_{k}}{\sqrt{N}}}\right]\left[\mathds{E}_{x^{\mu}_{i}}x^{\mu}_{i}\mathrm{e}^{-i\hat{h}^{\mu}\frac{x^{\mu}_{i}w^{*}_{i}}{\sqrt{N}}}\right]\left[\mathds{E}_{x^{\mu}_{j}}x^{\mu}_{j}\mathrm{e}^{-i\hat{h}^{\mu}\frac{x^{\mu}_{j}w^{*}_{j}}{\sqrt{N}}}\right]=\\ =\frac{w_{i}^{*}w_{j}^{*}}{N}\int Dh\,(h^{2}-1)\,\Theta\left(y^{\mu}\varphi(h)\right)=\frac{w_{i}^{*}w_{j}^{*}}{N}H_{2}(y^{\mu})\,$, (63)
@@ A.1 Annealed Computation
[…]
@@
$G_{S}^{(m)}(q,p_{1},\dots,p_{m})=\frac{1}{mN}\log{\frac{1}{2^{N}}\sum_{\boldsymbol{w}_{0}\cdots\boldsymbol{w}_{m}}\prod_{a<b}\delta\left(q-\frac{\boldsymbol{w}_{a}\cdot\boldsymbol{w}_{b}}{N}\right)\,\prod_{a}\delta\left(p_{a}-\frac{\boldsymbol{w}_{0}\cdot\boldsymbol{w}_{a}}{N}\right)}\,$, (70)
@@
$G_{S}^{(m)}(q,p_{1},\dots,p_{m})=\\ =\frac{1}{mN}\int\prod_{a}\frac{d\hat{p}_{a}}{2\pi/N}\prod_{a<b}\frac{d\hat{q}_{b}}{2\pi/N}\,\mathrm{e}^{-Nq\sum_{a<b}\hat{q}_{ab}-Np\sum_{a}\hat{p}_{a}}\left(\frac{1}{2}\sum_{w_{0}\cdots w_{m}}\mathrm{e}^{\sum_{a<b}\hat{q}_{ab}w_{a}w_{b}+w_{0}\sum_{a}\hat{p}_{a}w_{a}}\right)^{N}\,.$ (73)
@@
The last identity in Eq.˜77 follows noting that,
@@
$\frac{1}{2}\sum_{ab}q^{ab}\hat{h}^{a}\hat{h}^{b}=-\frac{q}{2}\left(\sum_{a}\hat{h}^{a}\right)^{2}-\frac{1-q}{2}\sum_{a}(\hat{h}^{a})^{2}\,.$ (81)